# Router-Tuning: A Simple and Effective Approach for Dynamic Depth

### v4

@@ 6 Ablation Studies
We conduct ablation studies to evaluate the generality and robustness of Router-Tuning across various settings, including different model architectures, the number of MoD layers, training datasets.
Table 4: Performance of Router-Tuning on Mixture of Experts, where we take Expert Drop He et al. (2024a) that statically skips the experts as the baseline of static dropping for comparison.
DeepSeek-MoE 1.00×1.$0$ 43.6 80.5 62.8 73.4 72.4 52.7 79.9 44.5 63.7
w/Drop $\bf 1.11\times$ 42.2 80.2 59.9 70.0 74.0 48.1 75.6 38.9 61.1
OLMoE 1.00×1.$0$ 45.6 80.1 53.7 71.2 74.7 54.5 79.4 52.5 64.0
w/Drop $\bf 1.13\times$ 34.0 66.6 51.6 59.3 67.6 39.0 40.5 42.7 50.2
@@
Figure 5: Ablation study on training samples.
@@
In the main experiments, we deployed half of the layers with MoD. In Figure 4, we further explore the effect of the number of MoD layers. Our results indicate that applying MoD to up to half of the attention layers still maintain comparable performance. However, when further increasing the number of MoD layers, performance starts to degrade. We attribute this decline to the transformation of important shallow layers, which negatively impacts overall performance Men et al. (2024); He et al. (2024a, b). Therefore, preserving the density of shallow layers while applying MoD to deeper layers ensures its effectiveness.
@@
7 Router-Tuning Extensions
In this section, we explore the extensions of Router-Tuning to broader architectural designs and training paradigms. Specifically, we demonstrate its applicability to mainstream Mixture-of-Experts (MoE) architectures, and examine its integration with LoRA-based fine-tuning to jointly improve efficiency and downstream performance.
Table 6: Effectiveness of Router-Tuning integrated with LoRA fine-tuning, compared to using Router-Tuning and LoRA independently. The integrated approach achieves comparable performance while preserving efficiency, highlighting the compatibility and complementary strengths of the two methods.
Model Method   OBQA PIQA   RTE WinoGrande   BoolQ ARC-C   HellaSwag MMLU   Avg.
Llama-3-8B Baseline 45.0 80.5 67.2 77.7 81.3 58.1 82.1 65.3 69.6
Router-Tuning 44.6 80.5 69.7 77.7 80.7 56.6 80.7 65.1 69.5
LoRA 46.6 82.0 68.0 77.9 83.9 61.8 81.6 65.9 71.0
   LoRA + Router-Tuning 47.2 82.2 67.4 77.8 83.9 61.5 81.7 65.8 70.9
Mistral-7B Baseline 44.4 82.2 68.2 79.1 82.2 60.6 83.2 62.4 70.3
Router-Tuning 44.2 81.9 68.5 78.6 81.7 60.4 82.5 61.8 70.0
LoRA 45.2 83.0 68.9 79.4 84.7 60.7 83.7 62.8 71.1
LoRA + Router-Tuning 45.7 83.1 68.7 79.3 84.3 60.9 83.4 62.9 71.2
@@ 7 Router-Tuning Extensions
The Expert’s redundancy in MoE has been widely demonstrated in recent works Lu et al. (2024); He et al. (2024a), e.g., the model still maintains comparable performance after removing certain layers. Therefore, we further extend Router-Tuning to MoE, where we take OLMoE Muennighoff et al. (2024) and DeepSeek-MoE Dai et al. (2024) as the backbones and equip each Expert network with Router-Tuning. Since these models deploy MoE at the token level, where each expert processes tokens from different sequences, we directly apply token-level Router-Tuning to these models. Here, we compare Router-Tuning with Expert Drop He et al. (2024a) that drops less important experts. For deployment, we configure an overall skipping ratio of 25%, achieved either by statically dropping 25% of layers or utilizing MoD on half of the MoE layers with 50% MoD capacity.
Figure 6: Expert-wise MoD capacity, visualizing the capacity utilization and selected frequency of individual experts within a specific layer.
In Table 4, while Expert Drop affects the performance significantly, Router-Tuning preserves more than 95% of the original performance, which demonstrates Router-Tuning’s effectiveness in the MoE architecture. In Figure 6, we further visualize the capacity utilization of individual experts within a specific layer, considering two perspectives for each expert: the frequency of expert selection (“Frequency”) and the MoD capacity allocated to each expert (“Capacity”).
Instead of removing a subset of experts (typically those being selected less frequently), Router-Tuning adopts a specialized and efficient utilization pattern, where high-frequency experts undergo greater sparsification compared to low-frequency experts. This approach retains the potential of all experts and fosters a more balanced utilization across the entire set of experts Fedus et al. (2022). Therefore, this approach maintains the potential of all experts and promotes a more balanced utilization across all experts.
Integration with LoRA Router-Tuning finetunes only a small subset of parameters, sharing similarities with parameter-efficient finetuning (PEFT) in terms of reduced parameter updates but differing in objectives: Router-Tuning focuses on achieving dynamic depth to enhance efficiency, whereas PEFT aims to adjust a small number of parameters to improve downstream task performance. Given that Router-Tuning is orthogonal to PEFT, we propose integrating Router-Tuning with LoRA finetuning Hu et al. (2021) to simultaneously enhance both task performance and efficiency. As the results shown in Table 6, jointly training routers and LoRA also preserves the performance of LoRA finetuning. Therefore, this joint training strategy serves as an advanced setting that further enhances the overall model capability.
@@ 8 Conclusion
In this work, we investigate the dynamic depth mechanism from both design and training perspectives. We propose Router-Tuning, which effectively enhances computational efficiency while preserving model performance. Furthermore, Router-Tuning implements dynamic depth by fine-tuning only a minimal number of parameters in just a few steps. Additionally, we explore Router-Tuning across a variety of modules and granularities to evaluate its effectiveness across a wide range of models and tasks. These advancements provide valuable insights and practical solutions for improving the efficiency of large language models.
@@
Despite the progress we have made, our work still has limitations. First, while we have advanced MoD with Router-Tuning, other sophisticated training methods may exist that could further improve performance, warranting future exploration. Second, due to computational resource constraints, our experiments have been limited to a small set of models and tasks. Expanding this approach to other models and a broader range of tasks would be highly valuable for understanding its full potential.

### v6

@@ 6 Ablation Studies
Figure 5: Ablation study on the impact of varying the number of MoD layers on overall model performance.
@@
In the main experiments, we deployed half of the layers with MoD. In Figure 5, we further explore the effect of the number of MoD layers. Our results indicate that applying MoD to up to half of the attention layers still maintains comparable performance. A similar trend is observed in Figure 6 for smaller models. However, when further increasing the number of MoD layers, performance starts to degrade. We attribute this decline to the transformation of important shallow layers, which negatively impacts overall performance Men et al. (2024); He et al. (2024a, b). Therefore, preserving the density of shallow layers while applying MoD to deeper layers ensures its effectiveness.
Figure 6: Performance of Router-Tuning on small models, where “R.T.” denotes Router-Tuning and the postfix “-nn” indicates that MoD is applied to $n$ layers.
@@
Figure 7: Effect of varying the number of training samples on performance.
@@
7 MoE and LoRA Integration
Table 7: Performance of Router-Tuning on Mixture of Experts, where we take Expert Drop He et al. (2024a) as the baseline of static dropping for comparison.
DeepSeek-MoE $1.00\times$ 43.6 80.5 62.8 73.4 72.4 52.7 79.9 44.5 63.7
w/Expert Drop $\bf 1.11\times$ 42.2 80.2 59.9 70.0 74.0 48.1 75.6 38.9 61.1
OLMoE $1.00\times$ 45.6 80.1 53.7 71.2 74.7 54.5 79.4 52.5 64.0
w/Expert Drop $\bf 1.13\times$ 34.0 66.6 51.6 59.3 67.6 39.0 40.5 42.7 50.2
In this section, we further explore the integration of Router-Tuning with other architectures and training techniques. First, we implement Router-Tuning on mainstream MoE architectures. Then, we combine Router-Tuning with LoRA fine-tuning to enhance both efficiency and performance.
Table 8: Effectiveness of Router-Tuning integrated with LoRA finetuning, compared to deploying Router-Tuning and LoRA separately.
Model   Method   SpeedUp    OBQA   PIQA   RTE   WinoGrande   BoolQ   ARC-C   HellaSwag   MMLU    Avg.
Llama-3-8B Baseline $1.00\times$ 45.0 80.5 67.2 77.7 81.3 58.1 82.1 65.3 69.6
R.T. $1.21\times$ 44.6 80.5 69.7 77.7 80.7 56.6 80.7 65.1 69.5
LoRA $1.00\times$ 46.6 82.0 68.0 77.9 83.9 61.8 81.6 65.9 71.0
LoRA + R.T. $1.21\times$ 47.2 82.2 67.4 77.8 83.9 61.5 81.7 65.8 70.9
Mistral-7B Baseline $1.00\times$ 44.4 82.2 68.2 79.1 82.2 60.6 83.2 62.4 70.3
R.T. $1.24\times$ 44.2 81.9 68.5 78.6 81.7 60.4 82.5 61.8 70.0
LoRA $1.00\times$ 45.2 83.0 68.9 79.4 84.7 60.7 83.7 62.8 71.1
  LoRA + R.T. $1.24\times$ 45.7 83.1 68.7 79.3 84.3 60.9 83.4 62.9 71.2
@@ 7 MoE and LoRA Integration
The Expert’s redundancy in MoE has been widely demonstrated in recent works Lu et al. (2024); He et al. (2024a), e.g., the model still maintains comparable performance after removing certain layers. Therefore, we further extend Router-Tuning to MoE, where we take OLMoE Muennighoff et al. (2024) and DeepSeek-MoE Dai et al. (2024) as the backbones and equip each Expert network with Router-Tuning. Since these models deploy MoE at the token level, we directly apply the token-level Router-Tuning to these models. Here, we compare Router-Tuning with Expert Drop He et al. (2024a) that statically drops less important experts. To ensure a fair comparison, we fix the overall skipping ratio of Router-Tuning to 25%, and drop the bottom 25% of experts globally by importance score in the Expert Drop baseline.
Figure 8: Normalized expert load before and after applying router-based filtering, denoted as “Assigned Tokens” and “Executed Tokens”, respectively. Expert load values are normalized by the mean number of assigned tokens.
In Table 7, instead of removing a subset of experts like Expert Drop, Router-Tuning maintains the potential of all experts, which contributes to a superior performance. In Figure 8, we further investigate how Router-Tuning affects the inference behavior of expert networks by visualizing the expert load within a specific layer from two perspectives: the number of tokens initially assigned to each expert (“Assigned Tokens”) and the number of tokens that pass the router and are actually executed (“Executed Tokens”). Router-tuning prompts the experts to skip less important tokens and significantly lower the load of overloaded experts, which alleviates the imbalanced distribution of token assignments. Consequently, this approach fosters a more balanced utilization across the entire set of experts Fedus et al. (2022); He et al. (2025) and thus enhances the efficiency.
Integration with LoRA
Router-Tuning enables dynamic depth to improve computational efficiency, whereas parameter-efficient fine-tuning (PEFT) methods aim to update a small subset of parameters to enhance downstream task performance. To examine whether Router-Tuning is complementary to PEFT, we propose jointly conducting Router-Tuning with LoRA fine-tuning Hu et al. (2021), targeting improvements in both efficiency and task performance. As shown in Table 8, this joint training strategy preserves the efficiency benefits of Router-Tuning while maintaining the performance gains achieved by LoRA. Together, the integration of Router-Tuning and LoRA offers a more advanced fine-tuning paradigm that further enhances overall model capability.
@@ 8 Conclusion
In this work, we investigate the dynamic depth mechanism from both design and training perspectives. We propose Router-Tuning, which effectively implements dynamic depth by fine-tuning only a minimal number of parameters in just a few steps. Additionally, we explore Router-Tuning across a variety of modules and granularities to evaluate its effectiveness across a wide range of models and tasks. These advancements provide valuable insights and practical solutions for deploying dynamic depth and enhancing the efficiency of large language models.
@@
Despite the progress achieved in this work, several limitations remain. First, while we have advanced MoD through Router-Tuning, other, potentially more sophisticated training strategies may further improve performance and merit future investigation. Second, due to computational resource constraints, our experiments were limited to a small set of models and tasks. Extending this approach to a broader range of architectures and applications would provide deeper insight into its generalizability and full potential.